OpenStudy (anonymous):

Factor and solve the equation: 9x^2 - 30x + 25 = 0

OpenStudy (anonymous):

9x^2-15x-15x+25=0 3x(3x-5)-5(3x-5)=0 (3x-5)(3x-5)=0

OpenStudy (anonymous):

(3x-5)^2=0 therfore ,x=5/3